Find the present value of $3,300 under each of the following rates and periods: (Round your final answer to the nearest penny.)

a. 8.9 percent compounded monthly for five years.

Present value $


b. 6.6 percent compounded quarterly for eight years.

Present value $


c. 4.3 percent compounded daily for four years.

Present value $


d. 5.7 percent compounded continuously for three years.

Present value $
Answer #1

Present value = Future value / (1+ (rate per period/no. of compounding per year)^(no. of periods*no. of compounding per year)

Answer a

Present value = 3300/(1+(.089/12))^(5*12)
= 3300/1.55792982687

= 2118

Answer b

Present value = 3300/(1+(.066/4))^(8*4)
= 3300/1.6882480438

= 1955

​​​​​​​Answer c

Present value = 3300/(1+(.043/365))^(4*365)
= 3300/1.18766580125

= 2779

​​​​​​​Answer d

Present value = Future value/e^(rt)
= 3300/e^(5.7*3)

= 3300/1.18649074902

= 2781